The scale of a map is 1:20,000.Calculate the area, in square centimeters, on the map of a forest reserve which covers 85kmē.
Answer by Boreal(15235)

1 cm is to 20000 cm (200 m)
as 5 cm is to 1 km (1000 m)
the map is 85 km^2
1 cm^2:4*10^8 cm^2
x cm^2=85(100000^2)=85*10^10 cm^2
Divide that and 21.25 *10^2 cm^2 on map
2125 cm^2.
Perhaps easier way
5 km x 17 km is 85 km^2
1 cm on map is 200 m, so 5 cm per km
This is 25 cm by 85 cm or 2125 cm^2